OpenStudy (anonymous):

Our teacher sent us to do some practice questions, but I don't get this one...could anyone help?

OpenStudy (michele_laino):

I think it is correct, so, we can apply the SAS criterion in order to say that the triangles ACD and ACB are congruent, and then AD is congruent to BC

OpenStudy (michele_laino):

oops.. we can apply the ASA criterion not SAS

OpenStudy (michele_laino):

sorry it was my typo!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

here, we can say that the ratio \(2:3\) holds for each pair of corresponding side, so we can write this: \[\frac{{AD}}{{EH}} = \frac{2}{3}\] therefore, we have: \[\frac{8}{{EH}} = \frac{2}{3}\]

OpenStudy (michele_laino):

sides*

OpenStudy (michele_laino):

please, what is \(x\)?

OpenStudy (michele_laino):

from my formula, above, I get: \[\huge EH = \frac{{3 \times 8}}{2} = ...?\]

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

no, since EH is the length measured on the photocopy, the original length is 8

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

here we have to draw both quadrilaterals using the system of coordinates \((x,y)\)

OpenStudy (michele_laino):

think that the have the same areas, since I got this drawing: 

OpenStudy (michele_laino):

they*

OpenStudy (michele_laino):

please wait I have made an error

OpenStudy (michele_laino):